OpenStudy (anonymous):

With a tail wind, a helicopter traveled 300mi in an hour and 40 min. the return trip against the same wind took 20 min longer. find the wind speed and also the air speed of the helicopter.

OpenStudy (anonymous):

You can use a equation to solve this question. Consider the speed of wind as x.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The equation is 1/3*(300+x)=1*(300-x), do you know why?

OpenStudy (anonymous):

no i got a system of equations.... I said p=air speed and w= wind speed 5/3(p+w)=300 2(p-w)=300

OpenStudy (anonymous):

My equation is easy to solve. Distance=time*speed, all right?

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

Is the speed of helicopter 300miles per hour?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The speed of the helicopter traveled with trial wind was 300+x, for wind was pushing it forward; the speed of the helicopter came back was 300-x, for wind was pushing it backward, isn't?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The distance the helicopter traveled can be represented by 1/3*(300+x) and 1*(300-x). So we get 1/3*(300+x)=1*(300-x).

OpenStudy (anonymous):

why 1/3?

OpenStudy (anonymous):

isnt it 5/3 though because it 1hr and 4o min

OpenStudy (anonymous):

Oh, I know. So the equation should be 5/3*(300+x)=2*(300-x), right this time?

OpenStudy (anonymous):

Thank for your reminding!

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

27.3?

OpenStudy (anonymous):

thats not the answer though